# Element2Vec: Build Chemical Element Representation from Text for Property Prediction

### v1

@@ 1 Introduction
With the recent advances of deep learning, the graph neural networks have emerged as powerful tools to implicitly learn element correlation for property prediction (LeCun et al., 2015; Shi et al., 2024; Omee et al., 2022; Pyzer-Knapp et al., 2022; Handoko and Made, 2025; Hu and Tian, 2023; Li, 2025), but may always require the data of properties used in input to be complete. In contrast, the data in materials science is of inherent and severe sparsity. Then, the large language models (LLMs) are finetuned for automatic chemical knowledge extraction from literature, followed by encoding in pre-defined formats (Van Herck et al., 2025; Xie et al., 2024; Jacobs et al., 2024), and direct material property prediction from chemical composition (Madika et al., 2025; Choudhary et al., 2022). Nevertheless, both approaches still rely on one-hot or handcrafted descriptors for elements, hindering the generalizability. Furthermore, issues like hallucination (Huang et al., 2025) and data distribution shift can compromise the reliability and utility of LLM further.
@@
Figure 1: Encode Element Representation from Text based on inspiration from Image Representation. (Left) Conventional image representation can be either a global embedding extracted from the whole image, or a set of local embeddings extracted from a sub-region that focuses on several key components Ma et al. (2024), such as tail and head. (Right) Analogously, given the tremendous amount of text in element description and its diverse application and aspects, we can also extract a single Global embedding from all text or following the definition in (Cardarelli, 2008) to extract Local embeddings highlighting different element attributes like atomic, chemical, thermal, and optical.
@@
In this work, we propose to employ LLM as an embedding extractor and construct a vector representation of chemical elements from free-form textual descriptions. For one thing, the materials are multifaceted entities with the description requiring a synthesis of physical, chemical, and structural attributes (Willatt et al., 2019). Then, the many properties, e.g., phase behavior such as gas and liquid, as well as the material design process, cannot be intuitively represented as mathematical descriptors. For another thing, the strong representation encoded with multifaceted characteristics may enable application of a simple algorithm for reliable and explainable prediction (Jordan and Mitchell, 2015; Yang et al., 2023). Thus, for each element, by integrating diverse information into a unified, context-rich representation (Jin et al., 2025), we may enable broader downstream applications.
@@
Specifically, we propose Element2Vecthat leverages the LLM to extract embeddings from text parsed from a Wikipedia webpage, and these embeddings can be directly compared in the same representation space. As illustrated in Fig. 1 (Left), motivated by the image encoding techniques that produce both a single embedding to describe entire image (globally) and a set of localized embeddings focusing on specific sub-regions, for each element, in addition to a Global embedding from the entire Wikipedia page to capture holistic information, we also design the streamline to extract a set of Local embeddings, where each is tailored to a specific attribute category, e.g., chemical and atomic, following the taxonomy defined in (Cardarelli, 2008). Without loss of generality, we assess properties framed as both classification and regression tasks. In particular, for regression under high data sparsity, we propose a test-time training method based on self-attention that recasts prediction as an imputation problem to enhance precision. With the contribution summarized as follows, we hope that our efforts could open new avenues for accelerating materials design (Jiang et al., 2025):
@@ 2 Related Work
AI for materials discovery. Deep learning graph networks such as CGCNN (LeCun et al., 2015) and MEGNet (Omee et al., 2022) have been applied for property prediction (Shi et al., 2024), while bond angles are further incorporated to improve predictions (Kaundinya et al., 2022; Choudhary and DeCost, 2021). Unsupervised natural language processing techniques have also uncovered chemical knowledge (Tshitoyan et al., 2019). With the recent success of language models, MatSciBERT (Gupta et al., 2022) is designed and tailored to materials text for information extraction, and LLMs like GPT-3 (Brown et al., 2020) are fine-tuned on chemistry tasks to predict molecular and materials properties competitively in low-data regimes (Van Herck et al., 2025; Xie et al., 2024; Jacobs et al., 2024).
@@
Chemical elements embedding. Early work used physical properties or compositional data to project elements into vector spaces that reflect periodic trends. For instance, methods based on self-organizing maps (Ward et al., 2018) and Magpie features (Lemes and Dal Pino, 2011) successfully grouped elements by similarity, while unsupervised approaches like Atom2Vec (Maaten and Hinton, 2008) and SkipAtom (Antunes et al., 2022) learned embeddings from compound structures that mirrored the periodicity of properties in an unsupervised manner. However, the dimensionality-reduction techniques (Maaten and Hinton, 2008; Jolliffe, 2011) primarily rediscover known relationships from existing data, lacking a mechanism for extrapolation beyond established knowledge. Nevertheless, all of the methods mentioned above depend on fixed or task-specific elemental encodings. Despite prior explorations with word embeddings (Yang et al., 2023) and materials knowledge graphs showing that language captures periodic trends and structure–property relations (Tshitoyan et al., 2019; Venugopal and Olivetti, 2024; Ye et al., 2024; Ma et al., 2021), the graph model still privileges geometric features but ignores language-informed semantics (Yan et al., 2024; Choudhary and DeCost, 2021; Reiser et al., 2022). Following the great success of large language models, recent efforts emphasize text mining, KG construction, or text-guided generation (Jiang et al., 2025; Das et al., 2025). Very recent “universal atomic embeddings” are trained from crystal data rather than textual descriptions (Jin et al., 2025). Still, the gap in constructing element-level embeddings from text still exists, which motivate our Element2Vec approach that learns broadly applicable element embeddings from text and evaluates them in material property prediction.
Large Language Models (LLM) have revolutionized text processing, enabling the flexible extraction of structured knowledge from unstructured text (Liu et al., 2023; Brown et al., 2020). Their capabilities in document segmentation, section identification, and intent classification allow for the precise capture of property-bearing statements from scientific literature (Xu et al., 2024; Dagdelen et al., 2024; Chen et al., 2025). Research in scientific NLP has developed domain-specialized LLMs to extract structured knowledge from text (Ling et al., 2023; Zhang et al., 2024; Gong et al., 2024). This has spurred the development of domain-specialized LLMs, like SciBERT (Gupta et al., 2024; Dagdelen et al., 2024), which demonstrate superior performance in scientific information extraction. Techniques such as prompting further enable complex tasks like classification and rationale generation without task-specific model fine-tuning. In our pipeline, we leverage these advancements to process Wikipedia text: an LLM segments it into attribute-specific units and filters noisy sentences, thus generating clean, structured input for subsequent embedding and prediction tasks.
Text-embedding models convert textual descriptions into fixed-size vectors that preserve semantic relationships (Wang et al., 2024). Early word embeddings like word2vec utilize word co-occurrence to discern word meanings and identify patterns (Mikolov et al., 2013). Nevertheless, they encounter difficulties in comprehending extended sentences and acquiring detailed knowledge about specific entities. Arora et al. (2017) proposed SIF as a simple yet strong baseline for sentence embeddings, and Logeswaran and Lee (2018) introduced QuickThoughts with a contrastive context-prediction objective. However, many text-only pipelines can be trained upon a sufficient amount of data and do not necessarily utilize structured knowledge, e.g., Wikipedia2Vec (Yamada et al., 2018), ERNIE (Sun et al., 2019), KEPLER (Wang et al., 2021). Recent large-scale embedding models leverage LLMs for data generation or as backbones (Wang et al., 2023; Li et al., 2023). In this work, we use Gemini Embedding, which is trained atop the Gemini LLM and explicitly designed to retain knowledge within fixed-size vectors (Lee et al., 2025).
@@
Figure 2: Element2Vec-Locals Embeddings Extraction. For each element, given the text parsed from the Wikipedia webpage, we use LLM to label each sentence about the attribute categories (Cardarelli, 2008) independently and summarize the full text to shorten the words via the prompts $\mathbf{p}_{c}$ and $\mathbf{p}_{g}$ respectively. For each subset, it is concatenated with the summary and fed into $f_{emb}$ for extraction.
@@ 3 Element2Vec: Chemical Elements Representation from Text
Element2Vec-Global. As an intuitive baseline, given the text corpus $\{\mathbf{s}_{i}\}_{i=1}^{N_{e}}$ of element $e$ that consists of $N_{e}$ sentence $\mathbf{s}_{i}$. The Global embedding is defined as a holistic representation of the entire corpus, which can be obtained as $f_{emb}(\{\mathbf{s}_{i}\}_{i=1}^{N_{e}})$ where $f_{emb}$ is implemented as Gemini embedding (Lee et al., 2025).
Despite the Global embedding’s aim for a comprehensive representation, a single embedding may struggle to explicitly represent all distinct attributes (Cardarelli, 2008) covered in the content. Meanwhile, the order at the attribute level depends on the sentence order, as the original sequential order of sentences preserved in the webpages also varies, the order-sensitive nature of $f_{emb}$ architecture can then introduce noise into the embeddings.
@@
In detail, following the concepts in (Cardarelli, 2008), we define eight attribute categories, i.e., “Mechanical (Physical)”, “Optical”, “Electrical and Magnetic”, “Thermal”, “Chemical”, “Atomic and Radiational”, “Applications”, and “Abundance”. Then, we classify each sentence $\mathbf{s}_{i}$ using a prompt $\mathbf{p}_{c}$ that instructs the LLM to assign it to an attribute, i.e., $\text{LLM}(concat(\mathbf{p}_{c},\mathbf{s}_{i}))$. Sentences tagged with the same attribute are automatically grouped into a distinct subset. Finally, by concatenating with a concise global context, $\text{LLM}(concat(\mathbf{p}_{s},\{\mathbf{s}_{i}\}_{i=1}^{N_{e}}))$, summarized via LLM by the prompt $\mathbf{p}$, we extract the corresponding Gemini embedding (Lee et al., 2025). The justification for including both components is provided by the ablation studies in Sec. 4.1.
@@ 4 Property Prediction by Element2Vec
We first evaluate the characteristics of the Global and Local embeddings via element classification Secs. 4.1 and 4.2. Subsequently, we apply these embeddings to property estimation, framing it as a regression problem and introducing and demonstrating the efficacy of the test-time training approach Secs. 4.3 and 4.4 even under conditions of high data sparsity.
@@ 4.3 Test-Time Training for Robust Property Value Imputation
Motivated by the fact that the properties of an element can be influenced by those of others, it is crucial to model inter-element correlations for accurate prediction. Consequently, we propose to employ a network $f_{sa}$ based on self-attention architecture (Vaswani et al., 2017) and design a corresponding test-time training pipeline (Ma et al., 2019) as shown in Fig. 5(a) as the training of network parameters and the property prediction of elements in $\mathcal{E}_{uk}$ can be done simultaneously.
In detail, we formulate the $\mathcal{E}_{k}\cup\mathcal{E}_{uk}$ as input sequence and fed into the self-attention layer, where the corresponding output of each input embedding is projected via a shared linear layer and used as the predicted property value. Then, for the network supervision, the objective is to only minimize the loss, e.g., mean squared error LeCun et al. (2015) of elements in $\mathcal{E}_{k}$ and no supervision will be applied for elements in $\mathcal{E}_{uk}$.
@@
For material research, since the atomic number, i.e., the indexes of the elements in the periodic table, is a critical factor in property prediction, the position encoding (Vaswani et al., 2017) should have been incorporated into our algorithm as inputs. However, this is unnecessary from our observation and we think the reason is because that the embedding extracted from the source text already encapsulate the corresponding information, thus rendering explicit encoding redundant.
@@ 4.4 Prediction of atomic properties
Following Sec. 4.2, to avoid confounding representation changes, all subsequent regression experiments therefore use the summary-prepended setting (local, front). Fig. 6 turns to a representative scalar property—the van der Waals (vdW) radius—both to motivate the task and to test whether text-derived embeddings carry element-level size information.
Fig. 6(a) shows the interaction–distance curve and the $R_{\mathrm{vdW}}$ relation used here, underscoring why $R_{\mathrm{vdW}}$ matter: they enter noncovalent contact a steric accessibility, packing, and force-field parameterization across chemistry, materials, and biomolecular modeling (Zefirov and Zorkii, 1989; Vemparala et al., 2004; Street and Mayo, 1998; Vanommeslaeghe et al., 2010). At the same time, $R_{\mathrm{vdW}}$ is difficult to determine and not uniquely defined; estimates depend on environment and methodology, often inferred from contact distributions, dispersion models, or polarizability trends rather than measured directly (Mantina et al., 2009; Tkatchenko and Scheffler, 2009; Rahm et al., 2016). This importance–difficulty trade-off motivates evaluating whether our embeddings can predict element-level $R_{\mathrm{vdW}}$.
@@
Followingly, given the limited dataset available for inferring atomic-level properties, we keep using the global embedding and then compare the performance of different predictive models in Fig. 7. First of all, we observe that test-time training consistently performs better than the traditional training approaches (both linear regression(LR) and MLP), particularly as the missing ratio increases. When comparing LR and MLP, when the labeled data is pretty limited, aggressively increasing the size of the network will actually make the prediction even worse. In contrast, our test-time training algorithm can adapt the model effectively to mitigate the impact of incomplete inputs, leading to substantially lower RMSE values across all missing ratios, especially at the higher ones.
@@ 5 Conclusion
Appendix A Appendix
@@ 0.2 He 1608 470 802 3083 2249 1971 1872 3565 3383
Figure 9: (a) The LR result of all Element2Vec-Locals embedding datasets with full attributes vector. (B) The LR result of all Element2Vec-Locals embedding datasets with only ARF attribute vector.

### v2

# Element2Vec: Learning Chemical Element Representation from Text for Property Prediction

@@ Abstract
keywords:
Large language model , Material science , Element encoding , Deep learning , Text mining , Data science
journal: ICLR 2026
@@ 1 Introduction
With the recent advances of deep learning, the graph neural networks have emerged as powerful tools to implicitly learn element correlation for property prediction [1, 2, 3, 4, 5, 6, 7], but may always require the data of properties used in input to be complete. In contrast, the data in materials science is of inherent and severe sparsity. Then, the large language models (LLMs) are finetuned for automatic chemical knowledge extraction from literature, followed by encoding in pre-defined formats [8, 9, 10], and direct material property prediction from chemical composition [11, 12]. Nevertheless, both approaches still rely on one-hot or handcrafted descriptors for elements, hindering the generalizability. Furthermore, issues like hallucination [13] and data distribution shift can compromise the reliability and utility of LLM further.
@@
Figure 1: Encode Element Representation from Text based on inspiration from Image Representation. (Left) Conventional image representation can be either a global embedding extracted from the whole image, or a set of local embeddings extracted from a sub-region that focuses on several key components [14], such as tail and head. (Right) Analogously, given the tremendous amount of text in element description and its diverse application and aspects, we can also extract a single Global embedding from all text or following the definition in [15] to extract Local embeddings highlighting different element attributes like atomic, chemical, thermal, and optical.
@@
In this work, we propose to employ LLM as an embedding extractor and construct a vector representation of chemical elements from free-form textual descriptions. For one thing, the materials are multifaceted entities with the description requiring a synthesis of physical, chemical, and structural attributes [16]. Then, the many properties, e.g., phase behavior such as gas and liquid, as well as the material design process, cannot be intuitively represented as mathematical descriptors. For another thing, the strong representation encoded with multifaceted characteristics may enable application of a simple algorithm for reliable and explainable prediction [17, 18]. Thus, for each element, by integrating diverse information into a unified, context-rich representation [19], we may enable broader downstream applications.
@@
Specifically, we propose Element2Vecthat leverages the LLM to extract embeddings from text parsed from a Wikipedia webpage, and these embeddings can be directly compared in the same representation space. As illustrated in Fig. 1 (Left), motivated by the image encoding techniques that produce both a single embedding to describe entire image (globally) and a set of localized embeddings focusing on specific sub-regions, for each element, in addition to a Global embedding from the entire Wikipedia page to capture holistic information, we also design the streamline to extract a set of Local embeddings, where each is tailored to a specific attribute category, e.g., chemical and atomic, following the taxonomy defined in [15]. Without loss of generality, we assess properties framed as both classification and regression tasks. In particular, for regression under high data sparsity, we propose a test-time training method based on self-attention that recasts prediction as an imputation problem to enhance precision. With the contribution summarized as follows, we hope that our efforts could open new avenues for accelerating materials design [20]:
@@ 2 Related work
AI for materials discovery. Deep learning graph networks such as CGCNN [1] and MEGNet [3] have been applied for property prediction [2], while bond angles are further incorporated to improve predictions [21, 22]. Unsupervised natural language processing techniques have also uncovered chemical knowledge [23]. With the recent success of language models, MatSciBERT [24] is designed and tailored to materials text for information extraction, and LLMs like GPT-3 [25] are fine-tuned on chemistry tasks to predict molecular and materials properties competitively in low-data regimes [8, 9, 10].
@@
Chemical elements embedding. Early work used physical properties or compositional data to project elements into vector spaces that reflect periodic trends. For instance, methods based on self-organizing maps [26] and Magpie features [27] successfully grouped elements by similarity, while unsupervised approaches like Atom2Vec [28] and SkipAtom [29] learned embeddings from compound structures that mirrored the periodicity of properties in an unsupervised manner. However, the dimensionality-reduction techniques [28, 30] primarily rediscover known relationships from existing data, lacking a mechanism for extrapolation beyond established knowledge. Nevertheless, all of the methods mentioned above depend on fixed or task-specific elemental encodings.
Despite prior explorations with word embeddings [18] and materials knowledge graphs showing that language captures periodic trends and structure–property relations [23, 31, 32, 33], the graph model still privileges geometric features but ignores language-informed semantics [34, 22, 35]. Following the great success of large language models, recent efforts emphasize text mining, KG construction, or text-guided generation [20, 36]. Very recent “universal atomic embeddings” are trained from crystal data rather than textual descriptions [19]. Still, the gap in constructing element-level embeddings from text still exists, which motivate our Element2Vec approach that learns broadly applicable element embeddings from text and evaluates them in material property prediction.
@@
Figure 2: Element2Vec-Locals Embeddings Extraction. For each element, given the text parsed from the Wikipedia webpage, we use LLM to label each sentence about the attribute categories [15] independently and summarize the full text to shorten the words via the prompts $\mathbf{p}_{c}$ and $\mathbf{p}_{g}$ respectively. For each subset, it is concatenated with the summary and fed into $f_{emb}$ for extraction.
Large Language Models (LLM) have revolutionized text processing, enabling the flexible extraction of structured knowledge from unstructured text [37, 25]. Their capabilities in document segmentation, section identification, and intent classification allow for the precise capture of property-bearing statements from scientific literature [38, 39, 40]. Research in scientific NLP has developed domain-specialized LLMs to extract structured knowledge from text [41, 42, 43].
This has spurred the development of domain-specialized LLMs, like SciBERT [44, 39], which demonstrate superior performance in scientific information extraction. Techniques such as prompting further enable complex tasks like classification and rationale generation without task-specific model fine-tuning. In our pipeline, we leverage these advancements to process Wikipedia text: an LLM segments it into attribute-specific units and filters noisy sentences, thus generating clean, structured input for subsequent embedding and prediction tasks.
Text-embedding models convert textual descriptions into fixed-size vectors that preserve semantic relationships [45]. Early word embeddings like word2vec utilize word co-occurrence to discern word meanings and identify patterns [46]. Nevertheless, they encounter difficulties in comprehending extended sentences and acquiring detailed knowledge about specific entities. arora2017simple [47] proposed SIF as a simple yet strong baseline for sentence embeddings, and logeswaran2018efficient [48] introduced QuickThoughts with a contrastive context-prediction objective. However, many text-only pipelines can be trained upon a sufficient amount of data and do not necessarily utilize structured knowledge, e.g., Wikipedia2Vec [49], ERNIE [50], KEPLER [51]. Recent large-scale embedding models leverage LLMs for data generation or as backbones [52, 53]. In this work, we use Gemini Embedding, which is trained atop the Gemini LLM and explicitly designed to retain knowledge within fixed-size vectors [54].
@@ 3 Element2Vec: Chemical Elements Representation from Text
Element2Vec-Global. As an intuitive baseline, given the text corpus $\{\mathbf{s}_{i}\}_{i=1}^{N_{e}}$ of element $e$ that consists of $N_{e}$ sentence $\mathbf{s}_{i}$. The Global embedding is defined as a holistic representation of the entire corpus, which can be obtained as $f_{emb}(\{\mathbf{s}_{i}\}_{i=1}^{N_{e}})$ where $f_{emb}$ is implemented as Gemini embedding [54].
Despite the Global embedding’s aim for a comprehensive representation, a single embedding may struggle to explicitly represent all distinct attributes [15] covered in the content. Meanwhile, the order at the attribute level depends on the sentence order, as the original sequential order of sentences preserved in the webpages also varies, the order-sensitive nature of $f_{emb}$ architecture can then introduce noise into the embeddings.
@@
In detail, following the concepts in [15], we define eight attribute categories, i.e., “Mechanical (Physical)”, “Optical”, “Electrical and Magnetic”, “Thermal”, “Chemical”, “Atomic and Radiational”, “Applications”, and “Abundance”. Then, we classify each sentence $\mathbf{s}_{i}$ using a prompt $\mathbf{p}_{c}$ that instructs the LLM to assign it to an attribute, i.e., $\text{LLM}(concat(\mathbf{p}_{c},\mathbf{s}_{i}))$. Sentences tagged with the same attribute are automatically grouped into a distinct subset. Finally, by concatenating with a concise global context, $\text{LLM}(concat(\mathbf{p}_{s},\{\mathbf{s}_{i}\}_{i=1}^{N_{e}}))$, summarized via LLM by the prompt $\mathbf{p}$, we extract the corresponding Gemini embedding [54]. The justification for including both components is provided by the ablation studies in Sec. 4.1.
@@ 4 Property Prediction by Element2Vec
We first evaluate the characteristics of the Global and Local embeddings via element classification Sec. 4.1 and  Sec. 4.2. Subsequently, we apply these embeddings to property estimation, framing it as a regression problem and introducing and demonstrating the efficacy of the test-time training approach Sec. 4.3 and Sec. 4.4 even under conditions of high data sparsity.
@@ 4.3 Test-Time Training for Robust Property Value Imputation
Motivated by the fact that the properties of an element can be influenced by those of others, it is crucial to model inter-element correlations for accurate prediction. Consequently, we propose to employ a network $f_{sa}$ based on self-attention architecture [55] and design a corresponding test-time training pipeline [56] as shown in Fig. 5(a) as the training of network parameters and the property prediction of elements in $\mathcal{E}_{uk}$ can be done simultaneously.
In detail, we formulate the $\mathcal{E}_{k}\cup\mathcal{E}_{uk}$ as input sequence and fed into the self-attention layer, where the corresponding output of each input embedding is projected via a shared linear layer and used as the predicted property value. Then, for the network supervision, the objective is to only minimize the loss, e.g., mean squared error [1] of elements in $\mathcal{E}_{k}$ and no supervision will be applied for elements in $\mathcal{E}_{uk}$.
@@
For material research, since the atomic number, i.e., the indexes of the elements in the periodic table, is a critical factor in property prediction, the position encoding [55] should have been incorporated into our algorithm as inputs. However, this is unnecessary from our observation and we think the reason is because that the embedding extracted from the source text already encapsulate the corresponding information, thus rendering explicit encoding redundant.
@@
Following Section 4.2, to avoid confounding representation changes, all subsequent regression experiments therefore use the summary-prepended setting (local, front). Fig. 6 turns to a representative scalar property—the van der Waals (vdW) radius—both to motivate the task and to test whether text-derived embeddings carry element-level size information.
Fig. 6(a) shows the interaction–distance curve and the $R_{\mathrm{vdW}}$ relation used here, underscoring why $R_{\mathrm{vdW}}$ matter: they enter noncovalent contact a steric accessibility, packing, and force-field parameterization across chemistry, materials, and biomolecular modeling [57, 58, 59, 60]. At the same time, $R_{\mathrm{vdW}}$ is difficult to determine and not uniquely defined; estimates depend on environment and methodology, often inferred from contact distributions, dispersion models, or polarizability trends rather than measured directly [61, 62, 63]. This importance–difficulty trade-off motivates evaluating whether our embeddings can predict element-level $R_{\mathrm{vdW}}$.
@@
Followingly, given the limited dataset available for inferring atomic-level properties, we keep using the global embedding and then compare the performance of different predictive models in Fig. 7.
First of all, we observe that test-time training consistently performs better than the traditional training approaches (both linear regression(LR) and MLP), particularly as the missing ratio increases. When comparing LR and MLP, when the labeled data is pretty limited, aggressively increasing the size of the network will actually make the prediction even worse. In contrast, our test-time training algorithm can adapt the model effectively to mitigate the impact of incomplete inputs, leading to substantially lower RMSE values across all missing ratios, especially at the higher ones.
@@ 5 Conclusion
Acknowledgments
This work was supported by the National Natural Science Foundation of China (52471178).
Appendix A Learning elemental properties by multi-layer perceptron
@@ 0.2 He 1608 470 802 3083 2249 1971 1872 3565 3383
Fig. 1: (a) The LR result of all Element2Vec-Locals embedding datasets with full attributes vector. (B) The LR result of all Element2Vec-Locals embedding datasets with only ARF attribute vector.